# Robust Training of Neural Networks at Arbitrary Precision and Sparsity

### v2

@@ Abstract
The discontinuous operations inherent in quantization and sparsification introduce a long-standing obstacle to backpropagation, particularly in ultra-low precision and sparse regimes. The standard Straight-Through Estimator (STE) is widely used to address this, but the well-understood mismatch between its quantization-aware forward pass and quantization-oblivious backward pass leads to unmanaged error that can corrupt the learning process. We solve this by introducing a denoising dequantization transform derived from a principled ridge regression objective. This transform makes the entire learning process aware of and robust to the quantization error that STE’s surrogate gradient bypasses, by creating an explicit, corrective gradient path. We extend this principle to sparsification by viewing it as a special form of quantization that maps insignificant values to zero. Our unified framework allows existing models to be trained at a wide spectrum of precisions and sparsity levels with off-the-shelf recipes, achieving stable training of fully binary (A1W1) and sparse sub-1-bit networks where other methods falter. This approach yields state-of-the-art results and provides a theoretically-grounded path to hyper-efficient neural networks.
@@ 1 Introduction
To deploy complex AI models on resource-constrained devices, techniques such as quantization and sparsification are essential. However, their non-differentiable nature poses a long-standing challenge for gradient-based training. For years, the community has relied on the Straight-Through Estimator (STE) (Bengio et al., 2013; Hubara et al., 2018), a surrogate gradient that has been instrumental in enabling Quantization-Aware Training (QAT). While STE has enabled progress, its use often leads to unpredictable and unstable convergence, particularly in demanding low-precision settings. This training instability is a well-documented obstacle; the estimator’s fragility is often concealed in large, over-parameterized models (Wang et al., 2023) but becomes apparent on smaller models that are more sensitive to quantization error, where STE-based methods frequently diverge (Figure 1(a)). This unreliability has necessitated a host of heuristic-based modifications and highlights the urgent need for a more principled and robust training method.
@@
Rather than layering on more empirical fixes, we propose a new QAT framework that addresses the source of this instability. The core mechanism of STE—approximating the derivative of the rounding function as an identity—creates a well-understood mismatch: the forward pass is affected by quantization error, while the backward pass is not. We posit that this lack of an explicit, corrective gradient path for the quantization error is a primary cause of the poor convergence that has long hindered ultra-low-bit QAT. Our approach is a non-straight-through method that, unlike STE, incorporates quantization error into the backward pass via a novel, data-dependent dequantization step. This method is simple to implement (Code Snippets  12), compatible with standard autograd systems, and provides a dual benefit: it creates a perturbation-resilient forward pass while ensuring a meaningful, error-aware gradient flows backward. This allows the network to learn to be robust to the quantization process itself, resulting in highly stable training and superior performance.
@@
Our framework’s robustness enables a more general approach to training efficient models. Whereas recent state-of-the-art methods often rely on bespoke recipe changes (Wang et al., 2023), architectural modifications, or complex, bit-specific recipe tuning (Liu et al., 2025), our method provides a universal “drop-in” solution that is effective across a wide range of precisions and standard architectures. It also allows us to unlock the full potential the theoretically superior but rarely used technique of affine quantization. The computational cost of a naive affine implementation has traditionally made it impractical; we overcome this barrier with a novel shortcut formula for affine quantized matrix multiplication that is both robust and efficient (Figure 1(b)).
@@
By treating sparsification as a form of quantization, our unified framework provides a stable foundation to explore the trade-offs between storage, energy, and quality. On a Gemma 1B model (Team et al., 2025), we map the storage-accuracy Pareto frontier (Figure 1(c)), revealing that asymmetric precision (e.g., 4-bit activations 1-bit weights (A4W1)) is optimal for storage. Furthermore, we analyze the trade-off between accuracy and computational efficiency by mapping an approximate energy-accuracy frontier (Figure 1(d)). For this, we use a hardware-agnostic cost metric that estimates the arithmetic computation effort Zhang et al. (2022), a dominant factor in the energy consumption of modern accelerators. This analysis reveals that structured sparsity can simultaneously reduce this approximate computational cost and improve accuracy. These findings support a powerful hypothesis: a larger model, when aggressively quantized, can outperform a smaller one at a superior efficiency footprint, pointing toward a new paradigm of hyper-efficient computing that echoes the remarkable efficiency of biological intelligence. Our key contributions are:
  • Pinpointing the “quantization-oblivious” backward pass of STE as a primary source of training instability.
@@
  • A simple, robust, and universal solution for training quantized and sparse networks based on a denoising dequantization transform derived from ridge regression, enabling, for the first time, stable A1W1 and sparse sub-1-bit training with standard, off-the-shelf recipes.
@@
(a) Training loss on Shakespeare Dataset
(b) Affine VS Linear Quantization
(c) Storage Efficiency Frontier
(d) Approximate Energy Efficiency Frontier
Figure 1: (a) Our method trains stably where standard QAT methods fail on small models. (b) It also unlocks the benefits of affine quantization, a gain STE fails to capture. This robustness provides a unified foundation to systematically map the storage (c) and energy (d) efficiency frontiers, revealing optimal trade-offs for state-of-the-art models.
@@ 2.1 Ghosts of Departed Quantities: The STE Blind Spot
The stability of Quantization-Aware Training (QAT) is critically dependent on its foundational algorithm: the Straight-Through Estimator (STE). A central element of our analysis is reformulating the standard quantization, $y=s\cdot\text{round}\left(\frac{\bm{x}}{s}\right)$ as an additive perturbation. By defining the rounding error $\delta=\text{round}(\frac{x}{s})-\frac{x}{s}$, a term that is critically detached and receives no gradient, the forward pass can be expressed as a simple addition: $y=s\cdot(\frac{x}{s}+\delta)=x+s\cdot\delta$.
@@ 2.3 The Promise of Bitwise Computation: A1W1 and Beyond
Solving these foundational issues enables a potential paradigm shift in computation: robustly training networks at extremely low precision, such as 1-bit for activations and weights (A1W1). At this level, expensive floating-point matrix multiplications can be replaced by highly efficient bitwise operations like XNOR and popcount (bit-counting). This creates a pathway for novel hardware architectures that are significantly simpler, faster, and more power-efficient. Furthermore, 1-bit networks offer a compelling link to biological computation, where neurons communicate not with high-precision values, but with discrete electrical impulses known as spikes (Hodgkin & Huxley, 1952; Dayan & Abbott, 2005). While Spiking Neural Networks (SNNs) aim to mimic this, their non-differentiable nature has made them difficult to train with gradient-based methods (Yamazaki et al., 2022; Lee et al., 2016; Tavanaei et al., 2019; Gallego et al., 2020). As this paper demonstrates, a robust algorithm for training 1-bit and even sub-1-bit networks helps bridge this long-standing gap between artificial and biological neural networks.
@@ 3.3 Stage 3: Dequantization with a Denoising Transform ($g$)
A core innovation of our work lies in the dequantization step which maps the quantized data back to the original floating-point range to approximate the unquantized data. While typical methods simply invert the scaling from Stage 1, our approach introduces a denoising dequantization transform, $g$, which is explicitly designed to solve the problem of the unmanaged quantization error, $\bm{\delta}$. We formulate this dequantization as a principled ridge regression problem, which has a powerful dual benefit: it creates a robust, differentiable path for the quantization error to participate in backpropagation, and the regularization factor $\lambda$ ensures numerical stability against low-variance data.
@@ 3.3.1 Formulation: General and Symmetric Transforms
The regularization parameter $\lambda$ acts as a “denoising” knob. As $\lambda\rightarrow\infty$, the scale $s_{g}\rightarrow 0$, and the dequantization collapses to $\overline{\bm{x}}$, forcing the transform to ignore the “noisy” quantized vector $\bm{q}$ and fall back to the most stable component of the original signal: its mean. This provides a mechanism to balance signal fidelity against noise suppression, which is critical for stabilizing the backward pass. A single value of $\lambda=0.01$ proved sufficient to ensure stable training across all our diverse experimental settings.
@@ 3.3.2 How This Solves the STE Blind Spot
2. The Backward Pass: The Gradient is Error-Aware. During backpropagation, the gradient with respect to the quantized vector q is computed via the chain rule: $\frac{dL}{d\bm{q}}=\frac{dL}{dg(\bm{q})}\frac{dg(\bm{q})}{d\bm{q}}$. Because the parameters of $g$ (its scale and offset) are calculated from the statistics of $\bm{q}$ (Eq. 3), its derivative is directly shaped by the values within $\bm{q}$. Since $\bm{q}$ contains the error $\bm{\delta}$, this local derivative $\frac{dg(\bm{q})}{d\bm{q}}$ becomes an explicit function of that error. By forcing the quantization error to participate in the backward pass, our transform provides the learning signal that STE discards, allowing preceding layers to adapt their weights and become robust to the error.
@@ 5.1 A Novel Shortcut Enabled by the $L_{2}$ Formulation
where variables with $X$ are column vectors (row-wise statistics) and variables with $W$ are column vectors (column-wise statistics) that are transposed where appropriate.
@@ 6 Experiments
We conduct a comprehensive set of experiments to validate the robustness and efficiency of our quantization-awre training framework. Our evaluation spans a range of model scales and architectures, from small-scale transformers to state-of-the-art Gemma LLMs (Team et al., 2025). A key finding is that while many standard quantization methods are benchmarked on large, overparameterized models where their flaws can be masked, these methods often falter on smaller models that have less redundancy. Our framework, in contrast, demonstrates superior stability and performance across all model scales.
@@ 6.1 Experimental Setup
To test the robustness of our framework, all hyperparameters including learning rate, optimizer settings (AdamW), and warmup/decay schedules were kept identical to those of the full-precision BF16 baseline models across all quantization experiments. We find that due to our robust formulation, the standard unquantized training recipe is not only sufficient but often optimal, eliminating the need for the complex, hand-tuned schedules frequently required by other methods. This adherence to a standard recipe provides a fair and challenging testbed for all quantization techniques evaluated.
All experiments employ fine-grained quantization, using either channel-wise or sub-channel quantization (SCQ) with a block size of $128$. SCQ is central to our strategy for handling the outlier dilemma. Instead of clipping outliers (which risks losing feature information), we use SCQ to localize their impact. The outlier is allowed to “blend” into the computation of the scaling factors (and bias) for its specific block, but is prevented from corrupting the parameters of the entire channel. While hardware manufacturers have introduced support for micro-blocks (e.g., of size 32) that make SCQ more hardware-friendly, we primarily evaluate our method using a larger block size of $128$ to thoroughly test its effectiveness, unless otherwise specified. We use the notation “AxWy” to denote x-bit activations and y-bit weights. Integer based quantizations are evaluated in this section. All transformer models are trained using the AdamW optimizer with a cosine learning rate schedule and warmup. The implementation frameworks are tailored to the experiment scale: small-scale experiments on nanoGPT models are conducted in PyTorch on NVIDIA A100 GPUs, whereas large-scale pre-training on Gemma LLMs is implemented in JAX and trained on up to 64 TPUs.
For our quantization-only experiments (1, 2, and 4-bit), we use affine quantization to effectively model asymmetric data distributions. In contrast, for all experiments involving ternary quantization (1.5-bit) or structured sparsity, we intentionally use the simpler linear quantization, which naturally produces representations favorable for hardware simplicity.
@@ 6.2 Core Method Validation on nanoGPT
We begin by validating our method’s core stability on small-scale GPT models from the nanoGPT repository (Karpathy, 2022). This low-resource setting highlights the fragility of standard methods. On the Shakespeare dataset, which contains approximately one million characters. The transformer model used has 6 layers, 6 attention heads, and 11 million parameters. This low-resource setting quickly highlights the fragility of standard methods. We apply channel-wise quantization on the model. As shown in Figure 1(a), our method, using its robust affine quantization, converges smoothly even when both activations and weights are quantized to 1-bit (A1W1). In a direct comparison, the affine version of STE exhibits significant instability at the same precision—an issue that could not be resolved by simply reducing the learning rate. For baseline comparisons, we utilize the publicly available symmetric implementations of BitNet (Wang et al., 2023) and ParetoQ (Liu et al., 2025), as they lack native support for affine quantization. Our results indicate that our method’s inherent stability enables it to effectively leverage the superior modeling capacity of affine quantization, a regime where other methods are either demonstrably unstable or lack native support.
@@ 6.3.1 The Storage Efficiency Frontier: Asymmetric Quantization is Key
Our first analysis examines the trade-off between model accuracy and the storage cost of the quantized weights, measured in effective bits per element (BPE). As illustrated in Figure 1(c), our method maps out a clear and compelling Pareto frontier, demonstrating a practical path to extreme model compression without significant performance degradation.
The central finding is that the optimal storage frontier is achieved not through a symmetric quantization scheme (e.g., A2W2), but through a distinctly asymmetric strategy: pairing robust 4-bit activations (A4) with aggressively quantized, ultra-low-bit weights (e.g. W1). This approach is effective because it allocates the precision budget intelligently, preserving activation precision to maintain information flow while leveraging the static nature of weights for extreme compression. By integrating this principle with structured sparsity, our framework can push weights into the sub-1-bit regime while preserving the model’s core capabilities.
@@
Our framework proposes a third, more direct approach. We also acknowledge outliers may be features and, like some coordinate rotation methods, we “blend” them; however, we blend their influence into the computation of the statistical dequantization parameters (Eq. 3) rather than the coordinates. Both outliers and normal samples participate in this computation, ensuring all samples contribute to the information flow. This is made robust through subchannel quantization (SCQ), which localizes this statistical blending. An outlier’s corrupting influence is contained within its own moderate-sized block, allowing its information to be preserved without distorting the quantization for the rest of the channel. Our results confirm that this strategy—local statistical blending via SCQ—defines a superior Pareto frontier, establishing a more direct and practical path to state-of-the-art storage efficiency than methods relying on complex transforms.

### v3

@@ Abstract
The discontinuous operations inherent in quantization and sparsification introduce a long-standing obstacle to backpropagation, particularly in ultra-low precision and sparse regimes. While the community has long viewed quantization as unfriendly to gradient descent due to its lack of smoothness, we pinpoint—for the first time—that the key issue is the absence of a proper gradient path that allows training to learn robustness to quantization noise. The standard Straight-Through Estimator (STE) exacerbates this with its well-understood mismatch: a quantization-aware forward pass but oblivious backward pass, leading to unmanaged error and instability. We solve this by explicitly modeling quantization as additive noise, making the full forward-backward path well-defined without heuristic gradient estimation. As one natural solution, we introduce a denoising dequantization transform derived from a principled ridge regression objective, creating an explicit, corrective gradient path that makes learning robust to the noise STE bypasses. We extend this to sparsification by treating it as a special form of quantization that zeros out small values. Our unified framework trains models at arbitrary precisions and sparsity levels with off-the-shelf recipes, enabling stable A1W1 and sub-1-bit networks where others falter. It yields state-of-the-art results, mapping efficiency frontiers for modern LLMs and providing a theoretically grounded path to hyper-efficient neural networks.
@@ 1 Introduction
To deploy complex AI models on resource-constrained devices, techniques such as quantization and sparsification are essential. However, their non-differentiable nature poses a long-standing challenge for gradient-based training. For years, the community has relied on the Straight-Through Estimator (STE) (Bengio et al., 2013; Hubara et al., 2018), a surrogate gradient that has been instrumental in enabling Quantization-Aware Training (QAT). While STE has enabled progress, its use often leads to unpredictable and unstable convergence, particularly in demanding low-precision settings. This training instability is a well-documented obstacle; the estimator’s fragility is often concealed in large, over-parameterized models (Wang et al., 2023) but becomes apparent on smaller models that are more sensitive to quantization error, where STE-based methods frequently diverge. Such instability has necessitated numerous heuristic modifications, including additional normalization (Zhang et al., 2022), learning rate adjustments Wang et al., 2023, optimizer changes (Liu et al., 2021a), and fine-tuning (McKinstry et al., 2018). These ad-hoc solutions underscore the urgent need for a more principled, robust training framework. Our work addresses this long-standing challenge with a theoretically grounded solution that does not rely on surrogate gradient estimation, instead deriving well-defined, differentiable gradients from a ridge regression objective.
@@
Rather than layering on more empirical fixes, we propose a new QAT framework that addresses the source of this instability. The core mechanism of STE—approximating the derivative of the rounding function as an identity—creates a well-understood mismatch: the forward pass is affected by quantization error, while the backward pass is not. This lack of an explicit, corrective gradient path for the quantization error causes poor convergence in ultra-low-bit QAT. Our non-straight-through method incorporates quantization error into the backward pass via a novel, data-dependent dequantization step (simple to implement; Code Snippets  12). It provides well-defined gradients without estimation, creating a perturbation-resilient forward pass and error-aware backward gradients.
@@
Our framework’s robustness enables a more general approach to training efficient models. Whereas recent state-of-the-art methods often rely on bespoke recipe changes (Wang et al., 2023), architectural modifications, or complex, bit-specific recipe tuning (Liu et al., 2025), our method provides a universal “drop-in” solution that is effective across a wide range of precisions and standard architectures. It also allows us to unlock the full potential of the theoretically superior but rarely used technique of affine quantization. The computational cost of a naive affine implementation has traditionally made it impractical; we overcome this barrier with a novel shortcut formula for affine quantized matrix multiplication that is both robust and efficient (Figure 1(b)).
@@
By treating sparsification as a form of quantization, our unified framework provides a stable foundation to explore the trade-offs between storage, energy, and quality. On a Gemma3 1B model (Team et al., 2025), we map the storage-accuracy Pareto frontier (Figure 2), revealing that asymmetric precision (e.g., 4-bit activations 1-bit weights (A4W1)) is optimal for storage. Furthermore, we analyze the trade-off between accuracy and computational efficiency by mapping an approximate energy-accuracy frontier (Figure 3). For this, we use a hardware-agnostic cost metric that estimates the arithmetic computation effort Zhang et al. (2022), a dominant factor in the energy consumption of modern accelerators. This analysis reveals that structured sparsity can simultaneously reduce this approximate computational cost and improve accuracy. These findings reinforce the viability of extremely low-bit precision for high-capacity models, suggesting that aggressive quantization is a key enabler for maximizing model performance within strict hardware constraints. Our key contributions are:
  • Pinpointing STE’s “quantization-oblivious” backward pass as the key instability source—revealing, for the first time, that proper gradients enable learning to handle quantization noise.
@@
  • A simple, robust denoising dequantization transform from ridge regression, enabling—for the first time—stable A1W1 and sub-1-bit training with standard recipes.
@@
(a) Training Stability on Shakespeare Dataset)
(b) Affine vs. Linear Quantization Performance
Figure 1: Training Stability and Quantization Robustness Analysis. (a) Comparison of training loss on the Shakespeare dataset with 1-bit weights and activations (A1W1). Standard STE and BitNet fail to stabilize in this extreme regime, exhibiting divergence or high loss. In contrast, our approach converges smoothly, matching the stability of higher-precision baselines. (b) Comparison of Linear vs. Affine quantization schemes at A1W1. Standard STE (right bars) fails to utilize the additional expressivity of affine parameters, showing no improvement over linear quantization. Our method (left bars) robustly learns the affine parameters (scale and bias), achieving a significant accuracy jump (see Table 1/Appendix A.1.2).
@@ 2.1 Ghosts of Departed Quantities: The STE Blind Spot
The stability of Quantization-Aware Training (QAT) is critically dependent on its foundational algorithm: the Straight-Through Estimator (STE). For over a decade, STE has been the de facto solution to the non-differentiable nature of quantization, but its reliance on surrogate gradient estimation has perpetuated training instability. A central element of our analysis is reformulating the standard quantization, $y=s\cdot\text{round}\left(\frac{\bm{x}}{s}\right)$ as an additive perturbation. By defining the rounding error $\delta=\text{round}(\frac{x}{s})-\frac{x}{s}$, a term that is critically detached and receives no gradient, the forward pass can be expressed as a simple addition: $y=s\cdot(\frac{x}{s}+\delta)=x+s\cdot\delta$.
@@ 2.3 The Promise of Bitwise Computation: A1W1 and Beyond
Solving these foundational issues facilitates the robust training of networks at extremely low precision, such as 1-bit for activations and weights (A1W1). At this level, expensive floating-point matrix multiplications can be replaced by highly efficient bitwise operations, specifically XNOR and popcount (bit-counting). This capability opens the door to hardware architectures that are significantly simpler, faster, and more power-efficient. Furthermore, while Spiking Neural Networks (SNNs) have long sought to achieve similar efficiency by mimicking discrete biological impulses (Hodgkin and Huxley, 1952; Dayan and Abbott, 2005), their non-differentiable nature poses significant training challenges (Yamazaki et al., 2022; Lee et al., 2016). As this paper demonstrates, our robust training algorithm for 1-bit networks offers a practical alternative, achieving comparable computational sparsity and efficiency within a standard gradient-based framework.
@@ 3.3 Stage 3: Dequantization with a Denoising Transform ($g$)
A core innovation of our work lies in the dequantization step which maps the quantized data back to the original floating-point range to approximate the unquantized data. While typical methods simply invert the scaling from Stage 1, our approach introduces a denoising dequantization transform, $g$, which is explicitly designed to solve the problem of the unmanaged quantization error, $\bm{\delta}$. Unlike heuristic gradient estimators, we formulate this dequantization as a principled ridge regression problem, yielding well-defined gradients that ensure numerical stability against low-variance data without any surrogate approximations.
@@ 3.3.1 Formulation: General and Symmetric Transforms
The regularization parameter $\lambda$ acts as a “denoising” knob. As $\lambda\rightarrow\infty$, the scale $s_{g}\rightarrow 0$, and the dequantization collapses to the mean: $\overline{\bm{x}}$, forcing the transform to ignore the “noisy” quantized vector $\bm{q}$ and fall back to the most stable component of the original signal: its mean. This provides a mechanism to balance signal fidelity against noise suppression, which is critical for stabilizing the backward pass. A single value of $\lambda=0.01$ proved sufficient to ensure stable training across all our diverse experimental settings.
@@ 3.3.2 How This Solves the STE Blind Spot
2. The Backward Pass: The Gradient is Error-Aware. During backpropagation, the gradient with respect to the quantized vector q is computed via the chain rule: $\frac{dL}{d\bm{q}}=\frac{dL}{dg(\bm{q})}\frac{dg(\bm{q})}{d\bm{q}}$. Because the parameters of $g$ (its scale and offset) are calculated from the statistics of $\bm{q}$ (Eq. 3), its derivative is directly shaped by the values within $\bm{q}$. Since $\bm{q}$ contains the error $\bm{\delta}$, this local derivative $\frac{dg(\bm{q})}{d\bm{q}}$ becomes an explicit function of that error (Sec. A.4). By forcing the quantization error to participate in the backward pass, our transform provides the learning signal that STE discards, allowing preceding layers to adapt their weights and become robust to the error.
@@ 5.1 A Novel Shortcut Enabled by the $L_{2}$ Formulation
where $\overline{\bm{x}}$ and $\overline{\bm{w}}$ denote the mean vectors, and $n$ represents the inner product dimension size.
@@
where $n$ is the inner dimension size, the bar notation $\overline{(\cdot)}$ denotes the mean, variables with $X$ are column vectors (row-wise statistics), and variables with $W$ are column vectors (column-wise statistics) that are transposed where appropriate.
@@ 6 Experiments
We conduct a comprehensive set of experiments to validate the robustness and efficiency of our quantization-aware training framework. Our evaluation spans a range of model scales and architectures, from small-scale transformers to state-of-the-art Gemma LLMs (Team et al., 2025). A key finding is that while many standard quantization methods are benchmarked on large, overparameterized models where their flaws can be masked, these methods often falter on smaller models that have less redundancy. Our framework, in contrast, demonstrates superior stability and performance across all model scales.
@@ 6.1 Experimental Setup
To test robustness, hyperparameters matched BF16 baselines. Our formulation eliminates need for tuned schedules. All use channel-wise or sub-channel quantization (SCQ, block 128) for fine-grained quantization. Notation: “AxWy” for x-bit activations, y-bit weights. For our quantization-only experiments (1, 2, and 4-bit), we use affine quantization to effectively model asymmetric data distributions. In contrast, for all experiments involving ternary quantization (1.5-bit) or structured sparsity, we intentionally use the simpler linear quantization, which naturally produces representations favorable for hardware simplicity. The implementation frameworks are tailored to the experiment scale: small-scale experiments on nanoGPT models are conducted in PyTorch on NVIDIA A100 GPUs, whereas large-scale pre-training on Gemma LLMs is implemented in JAX and trained on up to 64 TPUs.
@@ 6.2 Core Method Validation on nanoGPT
We begin by validating our method’s stability on small-scale GPT models from the nanoGPT repository (Karpathy, 2022). This low-resource setting effectively highlights the fragility of standard methods. We use the Shakespeare dataset, consisting of approximately one million characters, to train a transformer model with 6 layers, 6 attention heads, and 11 million parameters. We apply channel-wise quantization on the model. As shown in Figure 1(a), our method, using its robust affine quantization, converges smoothly even when both activations and weights are quantized to 1-bit (A1W1). In a direct comparison, the affine version of STE exhibits significant instability at the same precision—an issue that could not be resolved by simply reducing the learning rate. These results validate our theoretically grounded solution, which succeeds without gradient estimation where long-standing methods fail. For baseline comparisons, we utilize the publicly available symmetric implementations of BitNet (Wang et al., 2023) and ParetoQ (Liu et al., 2025), as they lack native support for affine quantization. Our results indicate that our method’s inherent stability enables it to effectively leverage the superior modeling capacity of affine quantization, a regime where other methods are either demonstrably unstable or lack native support.
@@ 6.3.1 The Storage Efficiency Frontier: Asymmetric Quantization is Key
Our first analysis examines the trade-off between model accuracy and the storage cost of the quantized weights, measured in effective bits per element (BPE). As illustrated in Figure 2, our method maps out a clear and compelling Pareto frontier, demonstrating a practical path to extreme model compression without significant performance degradation.
The central finding is that the optimal storage frontier is achieved not through a symmetric quantization scheme (e.g., A2W2), but through a distinctly asymmetric strategy: pairing robust 4-bit activations (A4) with aggressively quantized, ultra-low-bit weights (e.g., W1). This approach is effective because it allocates the precision budget intelligently, preserving activation precision to maintain information flow while leveraging the static nature of weights for extreme compression. By integrating this principle with structured sparsity, our framework can push weights into the sub-1-bit regime while preserving the model’s core capabilities.
@@
Our framework proposes a third, more direct approach. We also acknowledge outliers may be features and, like some coordinate rotation methods, we “blend” them; however, we blend their influence into the computation of the statistical dequantization parameters (Eq. 3) rather than the coordinates. Both outliers and normal samples participate in this computation, ensuring all samples contribute to the information flow. This is made robust through subchannel quantization (SCQ), which localizes this statistical blending. An outlier’s corrupting influence is contained within its own moderate-sized block, allowing its information to be preserved without distorting the quantization for the rest of the channel. Our results confirm that this strategy—local statistical blending via SCQ—defines a superior Pareto frontier, establishing a more direct and practical path to state-of-the-art storage efficiency than methods relying on complex transforms like Hadamard.
Figure 2: Storage Efficiency Frontiers: We map the trade-off between validation accuracy and effective bits-per-element (BPE). The frontier reveals that asymmetric quantization (e.g., A4W1) provides a superior storage-accuracy trade-off compared to symmetric settings (e.g., A2W2), as it preserves activation information while aggressively compressing static weights.
Figure 3: Approximate Energy Efficiency Frontier: Estimating compute cost (Activation Bits $\times$ Weight Bits $\times$ Sparsity). The results demonstrate a synergy between our estimator and structured sparsity: the (1:4 and 2:4) structured sparse A4W1 models reduce the compute cost of the dense equivalent while maintaining high accuracy, defining the efficiency frontier.